OpenStudy (sissyedgar):

What is the slope-intercept form of the function that contains the points (3, 1) and (2, 9)? y = ___ x + ___ Thankiez!!!! ღ

OpenStudy (sissyedgar):

@TheSmartOne @StudyGurl14

TheSmartOne (thesmartone):

We first need to find the slope! :) \(\sf\Large Slope = \frac{\color{blue}{y_2}-\color{red}{y_1}}{\color{green}{x_2}-\color{purple}{x_1}}\) Where the two points are in the form: \(\sf\Large (\color{purple}{x_1},\color{red}{y_1})~and~(\color{green}{x_2},\color{blue}{y_2})\)

OpenStudy (sissyedgar):

Sorry internet was laggy! ( 3,1 ) ღ ( 2,9 )

TheSmartOne (thesmartone):

so those are the two points, now we plug them in. I even added colors to make it easier! :)

OpenStudy (sissyedgar):

So.... y = 8x + -1 Right? ღ

TheSmartOne (thesmartone):

slope is -8 not 8, so re-do the y-intercept :)

OpenStudy (sissyedgar):

So, y = -8x + -1

TheSmartOne (thesmartone):

nope y = -8x + b plug in x = 3, and y = 1 and solve for b

OpenStudy (sissyedgar):

-2

TheSmartOne (thesmartone):

how?

OpenStudy (sissyedgar):

3-1=-2

TheSmartOne (thesmartone):

your ignoring the -8

TheSmartOne (thesmartone):

1 = -8(3) + b

OpenStudy (sissyedgar):

1 = -8(3) + -2?

TheSmartOne (thesmartone):

nope, w =on't work what is 8 * 3 = ?

OpenStudy (sissyedgar):

8*3=24

TheSmartOne (thesmartone):

and so -8 * 3 = -24 and 1 = -24 + b

OpenStudy (sissyedgar):

is the b -2?

TheSmartOne (thesmartone):

noooooooooooo 1 = -24 + b add 24 on both sides what is 24 + 1 = ?

OpenStudy (sissyedgar):

25

OpenStudy (sissyedgar):

24+1=25 -24+24=0

TheSmartOne (thesmartone):

so what does b = ?

OpenStudy (sissyedgar):

25!

TheSmartOne (thesmartone):

Correct! ღ Great job! ღ

OpenStudy (sissyedgar):

My Heart! ღ xP

OpenStudy (sissyedgar):

Quick Question tho... Is 25 what comes after the 'x+'? xP

OpenStudy (sissyedgar):

@TheSmartOne

TheSmartOne (thesmartone):

yes so y = -8x + 25

OpenStudy (sissyedgar):

oooh, ok! Thankiez!!! ღ